OpenStudy (anonymous):

Multiply and simplify: x squared minus 4 over x squared minus 1 times x + 1 over x + 2

OpenStudy (eyust707):

do you know the trick when you have a "difference of a perfect square"

OpenStudy (anonymous):

no

OpenStudy (eyust707):



OpenStudy (eyust707):

notice that x^2 is a perfect square because x * x = x^2

OpenStudy (eyust707):

and 4 is a perfect square because 2 * 2 = 4

OpenStudy (eyust707):

and it is a difference because there is a minus sign between them

OpenStudy (anonymous):

oh ok

OpenStudy (eyust707):

when we satisfy those three thing theres this cool trick to factor it that always works

OpenStudy (eyust707):